Find the missing resistance R in the given circuit. Assume the current i=2.72 amps.

AOK ट-४ US 19
Answer #1

equivalent resistance of circuit is Req - GXR15 so ex P) + 5 72727 201 (6+R/ 6XRX2-12 + 13.6 = 20 6+R = 16-32.B = 64 6+R 16.3